OpenStudy (anonymous):

the odometer of a car reads 2000km at the start of the trip and 2400km at the end of the trip.if the trip took 8 hr,calculate the average speed of the car in km /h and m /s

OpenStudy (anonymous):

@ash2326 @apoorvk @A.Avinash_Goutham @nphuongsun93 @lgbasallote @myininaya @mathslover

OpenStudy (anonymous):

please help

mathslover (mathslover):

what did u try ? where r u facing problem ?

OpenStudy (anonymous):

please help dudes

mathslover (mathslover):

Readigs at odometer before the trip = 2000km Readigs at odometer after the trip = 2400km Distance travelled = 2400-2000=400 km right ?

OpenStudy (anonymous):

its set of questions i found in book and i believe that u all will help me

OpenStudy (anonymous):

ya then

mathslover (mathslover):

time take = 8 hrs Average speed in Km/h=400/8=50 So Average speed in Km/h=50 km/h

OpenStudy (anonymous):

thanks got it please check my answer okay guys

mathslover (mathslover):

and btw for transparency !!! http://www.meritnation.com/discuss/question/1008352 This is the link

OpenStudy (anonymous):